# Impact of Small-Scale Gravity Waves on Tracer Transport

### v1

@@ 2.2 Leading-order gravity wave tracer flux convergence
In this section, we describe how to calculate the leading-order gravity wave tracer flux convergence, $\mathcal{Q}^{(0)}$, in Equation 24 using the wave action density and wavenumbers predicted by a gravity wave model.
We aim to describe $\mathcal{Q}^{(0)}$ in terms of the wave action density, $\mathcal{A}_{\alpha}=E_{w,\alpha}/\hat{\omega}_{\alpha}$, where $E_{w,\alpha}$ is the wave energy given by Achatz et al. [1] as
$E_{w,\alpha}=\frac{\overline{\rho}}{2}\left(\frac{\left|\hat{\mathbf{U}}_{\alpha 1}^{(0)}\right|^{2}}{2}+\varepsilon^{5-\delta}\frac{\left|\hat{W}_{\alpha 1}^{(0)}\right|^{2}}{2}+\frac{1}{N_{0}^{2}}\frac{\left|\hat{B}_{\alpha 1}^{(0)}\right|^{2}}{2}\right)=\overline{\rho}\frac{\left|\hat{B}_{\alpha 1}^{(0)}\right|^{2}}{2N_{0}^{2}}\frac{N_{0}^{2}k_{h,\alpha}^{2}+f_{0}^{2}m_{\alpha}^{2}}{N_{0}^{2}k_{h,\alpha}^{2}}\;,$ (26)
$\hat{B}_{\alpha 1}^{(0)}=\hat{\Theta}_{\alpha 1}^{(0)}/\overline{\underline{\Theta}}$ is the leading-order buoyancy amplitude, and the horizontal wavenumber $k_{h,\alpha}$ is defined as $k_{h,\alpha}=\sqrt{k_{\alpha}^{2}+l_{\alpha}^{2}}$. The dispersion relation for internal gravity waves is given by
$\hat{\omega}_{\alpha}^{2}=\frac{N_{0}^{2}k_{h,\alpha}^{2}+f_{0}^{2}m_{\alpha}^{2}}{\varepsilon^{5-\delta}k_{h,\alpha}^{2}+m_{\alpha}^{2}}\;.$ (27)
Therefore, we must express the leading-order tracer and wind amplitudes, $\hat{\Psi}_{\alpha 1}^{(\gamma+1)}$ and $\hat{\mathbf{V}}_{\alpha 1}^{(0)}$, respectively, in terms of the buoyancy amplitude $\hat{B}_{\alpha 1}^{(0)}$. The tracer amplitude is coupled to the wind amplitude via Equation 17. The wind amplitude, in turn, is couple to the buoyancy amplitude through the polarization relations [3, 1]
@@
$M_{\alpha 1}\mathbf{Z}_{\alpha 1}^{(0)}=0\;,$ (31)
where the matrix $M_{\alpha 1}$ is given by
the components of the wave numbers are denoted by $\mathbf{k}_{\alpha}=(k_{\alpha},l_{\alpha},m_{\alpha})^{\mathrm{T}}$, and the vector $\mathbf{Z}_{\alpha 1}^{(0)}$ contains the wave amplitudes
$\mathbf{Z}_{\alpha 1}^{(0)}=\left(\hat{U}_{\alpha 1}^{(0)}\>,\>\hat{V}_{\alpha 1}^{(0)}\>,\>\hat{W}_{\alpha 1}^{(0)}\>,\>\hat{B}_{\alpha 1}^{(0)}/N_{0}\>,\>\frac{c_{p}}{R}\overline{\underline{\Theta}}\hat{\Pi}_{\alpha 1}^{(0)}\right)^{\mathrm{T}}\;.$ (33)
@@
By inserting Equations 17, 28, 29, 27 and 26 into Equation 24, we obtain an expression for $\mathcal{Q}^{(0)}$ in terms of the wave action density:
$\mathcal{Q}^{(0)}=-\frac{\varepsilon^{(2\gamma+1)}}{\overline{\rho}}\sum_{\alpha=1}^{A}\nabla_{\mathbf{X}}\cdot\left(\frac{f_{0}}{\hat{\omega}_{\alpha}}\frac{m_{\alpha}}{\varepsilon^{5-\delta}k_{h,\alpha}^{2}+m_{\alpha}^{2}}\mathcal{A}_{\alpha}\mathbf{k}_{\alpha}\times\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\right)\;.$ (34)
@@ 2.3 Next-order gravity wave tracer flux convergence
Similar to the previous section, we now describe how the next-order flux convergence, $\mathcal{Q}^{(1)}$, can be determined from the wave action density.
The next-order flux convergence, $\mathcal{Q}^{(1)}$, is given in Equation 25. Since $\mathcal{A}_{\alpha}\propto\left|\hat{B}_{\alpha 1}^{(0)}\right|^{2}$, both the leading-order and next-order tracer and wind amplitudes must be expressed in terms of the buoyancy amplitude. As described in the previous section, the leading-order amplitudes are coupled to $\hat{B}_{\alpha 1}^{(0)}$ via Equations 17, 28 and 29. The next-order tracer amplitudes can be obtained from the next-order wind amplitudes and the leading-order tracer amplitudes, as shown in Equation 18. Therefore, the remaining task is to determine $\hat{\mathbf{V}}_{\alpha 1}^{(1)}$ in terms of $\hat{B}_{\alpha 1}^{(0)}$. Similar to the leading-order basic wave amplitudes, the next-order wind, buoyancy and Exner pressure amplitudes $\hat{\mathbf{V}}_{\alpha 1}^{(1)}$, $\hat{B}_{\alpha 1}^{(1)}=\hat{\Theta}_{\alpha 1}^{(1)}/\overline{\underline{\Theta}}$, and $\hat{\Pi}_{\alpha 1}^{(1)}$, respectively, satisfy a matrix equation
$M_{\alpha 1}\mathbf{Z}_{\alpha 1}^{(1)}=\mathbf{R}_{\alpha 1}^{(1)}\;,$ (35)
where $M_{\alpha 1}$ given in Equation 32,
@@
The components of $\mathbf{R}_{\alpha 1}^{(1)}$ are
with $\mathbf{R}_{\mathbf{u},\alpha 1}^{(1)}=\left(R_{u,\alpha 1}^{(1)}\>,\>R_{v,\alpha 1}^{(1)}\right)^{\mathrm{T}}$. These right-hand sides follow Achatz et al. [3]. However, in contrast to that work, we have included higher-order terms from the case $\delta=1$ to facilitate a unified treatment of the two stratification cases. Likewise, we retain the full expansion of the advection large-scale winds. Using the polarization relations from Equations 28, 29 and 30, all right-hand sides can be determined form the spatial and temporal dependence of the wavenumber and leading-order buoyancy amplitude $\hat{B}_{\alpha 1}^{(0)}$. Once these terms are known, the next-order wave amplitudes $\mathbf{Z}_{\alpha 1}^{(1)}$ can be obtained by solving Equation 35. It is important to note that the next-order amplitude vector $\mathbf{Z}_{\alpha 1}^{(1)}$ does not belong to the null space of $M_{\alpha 1}$, as the latter is assumed to only contribute to the basic-wave amplitudes $\mathbf{Z}_{\alpha 1}^{(0)}$.
@@
$\hat{B}_{\alpha 1}^{(0)}=\left|\hat{B}_{\alpha 1}^{(0)}\right|e^{i\varLambda_{\alpha}}\;,$ (42)
where $\varLambda_{\alpha}$ represents the large-scale wave amplitude. The absolute magnitude can be obtained from the wave energy $E_{w,\alpha}$ via Equation 26. While $E_{w,\alpha}$ is directly available from the gravity wave model, the large-scale phase must be predicted from the prognostic equation
$\begin{split}\left(\partial_{\mathrm{T}}+\mathbf{c}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\varLambda_{\alpha}=&-\frac{f_{0}m_{\alpha}}{2\left|\mathbf{k}_{\alpha}\right|^{2}\hat{\omega}_{\alpha}}\left(\mathbf{k}_{\alpha}\times\nabla_{\mathbf{X}}\right)\left\langle{\underline{\mathbf{U}}}\right\rangle\\ &+\frac{N_{0}^{2}}{2\left(N_{0}^{2}-\varepsilon^{5-\delta}f_{0}^{2}\right)m_{\alpha}}\left[\varepsilon^{\delta}\frac{c_{p}}{R}\overline{\underline{\Theta}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Pi}}\right\rangle+\frac{1}{N_{0}^{2}}\frac{1}{\overline{\underline{\Theta}}}\left(\hat{\mathbf{c}}_{g,\alpha}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Theta}}\right\rangle\right]\\ &-\frac{f_{0}m_{\alpha}^{2}}{\left|\mathbf{k}_{\alpha}\right|^{4}\hat{\omega}_{\alpha}}\left[\left(k_{\alpha}\partial_{\mathrm{Y}}-l_{\alpha}\partial_{\mathrm{X}}\right)\hat{\omega}_{\alpha}+l_{\alpha}\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)k_{\alpha}-k_{\alpha}\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)l_{\alpha}\right]\;,\end{split}$ (43)
where $\mathbf{c}_{g,\alpha}=\nabla_{\mathbf{k}}\omega_{\alpha}$ is the group velocity and $\hat{\mathbf{c}}_{g,\alpha}=\nabla_{\mathbf{k}}\hat{\omega}_{\alpha}$ is the intrinsic group velocity. We derive the prognostic equation by taking the imaginary part of the solvability condition, as outlined in Achatz et al. [2]:
$0=\mathbf{Z}_{\alpha 1}^{(0)^{\dagger}}\mathbf{R}_{\alpha 1}^{(1)}\;.$ (44)
In summary, $\hat{B}_{\alpha 1}^{(0)}$ in Equation 42 can be calculated using Equations 26 and 43. The leading-order tracer and wind amplitudes can then be determined from Equation 17 and Equations 28 and 29, respectively. Next, by solving Equation 35 for $\mathbf{Z}_{\alpha 1}^{(1)}$ and using Equation 18, the next-order tracer and wind amplitudes can be computed. Finally, the resulting $\hat{\Psi}_{\alpha 1}^{(\gamma+1)}$, can be substituted into Equation 25 to obtain $\mathcal{Q}^{(1)}$.
@@ 3 Summary of equations in dimensional form
$\left[\left\langle{\underline{\mathbf{U}}}\right\rangle,\>\left\langle{\underline{W}}\right\rangle,\>\left\langle{\underline{\Psi}}\right\rangle,\>\left\langle{\underline{\Theta}}\right\rangle,\>\left\langle{\underline{\Pi}}\right\rangle\right]$ $\rightarrow\left[\frac{\left\langle{\mathbf{u}}\right\rangle}{\widetilde{U}},\>\frac{\left\langle{w}\right\rangle}{\varepsilon\widetilde{W}},\>\frac{\left\langle{\psi}\right\rangle}{\widetilde{\Psi}},\>\frac{\left\langle{\theta}\right\rangle}{\varepsilon^{\delta+1}T_{00}},\>\frac{\left\langle{\pi}\right\rangle}{\varepsilon^{\delta+1}}\right]\;.$ (51)
@@
$\hat{\omega}_{\alpha}^{2}=\frac{N^{2}k_{h,\alpha}^{2}+f^{2}m_{\alpha}^{2}}{k_{h,\alpha}^{2}+m_{\alpha}^{2}}\;,$ (52)
@@
$\left(\hat{\mathbf{U}}_{\alpha 1}^{(0)}\>,\>\hat{W}_{\alpha 1}^{(0)}\>,\>\hat{B}_{\alpha 1}^{(0)}\>,\>\hat{\Pi}_{\alpha 1}^{(0)}\>,\>\hat{\Psi}_{\alpha 1}^{(\gamma+1)}\right)\rightarrow\left(\frac{\mathbf{u}_{\alpha}^{\prime}}{\widetilde{U}}\>,\>\frac{w_{\alpha}^{\prime}}{\widetilde{W}}\>,\>\frac{\theta_{\alpha}^{\prime}}{\varepsilon^{\delta+1}\overline{\theta}}\>,\>\frac{\pi_{\alpha}^{\prime}}{\varepsilon^{\delta+2}}\>,\>\frac{\psi_{\alpha}^{\prime}}{\varepsilon\widetilde{\Psi}}\right)\frac{1}{\varepsilon^{\gamma}}\;,$ (53)
@@
$\left(\hat{\mathbf{U}}_{\alpha 1}^{(1)}\>,\>\hat{W}_{\alpha 1}^{(1)}\>,\>\hat{B}_{\alpha 1}^{(1)}\>,\>\hat{\Pi}_{\alpha 1}^{(1)}\>,\>\hat{\Psi}_{\alpha 1}^{(\gamma+2)}\right)\rightarrow\left(\frac{\mathbf{u}_{\alpha}^{\prime\prime}}{\widetilde{U}}\>,\>\frac{w_{\alpha}^{\prime\prime}}{\widetilde{W}}\>,\>\frac{\theta_{\alpha}^{\prime\prime}}{\varepsilon^{\delta+1}\overline{\theta}}\>,\>\frac{\pi_{\alpha}^{\prime\prime}}{\varepsilon^{\delta+2}}\>,\>\frac{\psi_{\alpha}^{\prime\prime}}{\varepsilon\widetilde{\Psi}}\right)\frac{1}{\varepsilon^{\gamma+1}}\;,$ (59)
and the components of $\mathbf{R}_{\alpha}$ are
@@
Re-dimensionalized, the gravity wave-induced tracer Stokes drift $\mathcal{Q}^{(0)}$ is
while the next-order gravity wave-impact
$\begin{split}\mathcal{Q}^{(1)}=-\frac{1}{2\overline{\rho}\overline{\theta}}\sum_{\alpha=1}^{A}&\nabla\cdot\mathfrak{R}\left[\overline{\rho}\overline{\theta}\left(\mathbf{v}_{\alpha}^{\prime}\psi_{\alpha}^{\prime\prime*}+\mathbf{v}_{\alpha}^{\prime\prime}\psi_{\alpha}^{\prime*}\right)\right]\end{split}$ (66)
is determined through a sequential process:

### v2

@@ 2.2 Leading-order gravity wave tracer flux convergence
In this section, we describe how to calculate the leading-order gravity wave tracer flux convergence, $\mathcal{Q}^{(0)}$, in Equation˜25 using the wave action density and wavenumbers predicted by a gravity wave model.
We aim to describe $\mathcal{Q}^{(0)}$ in terms of the wave action density, $\mathcal{A}=E_{w}/\hat{\omega}$, where $E_{w,}$ is the wave energy given by Achatz et al. [2] as
$E_{w}=\frac{\overline{\rho}}{2}\left(\frac{\left|\hat{\mathbf{U}}^{(0)}\right|^{2}}{2}+\varepsilon^{4}\frac{\left|\hat{W}^{(0)}\right|^{2}}{2}+\frac{1}{N_{0}^{2}}\frac{\left|\hat{B}^{(0)}\right|^{2}}{2}\right)=\overline{\rho}\frac{\left|\hat{B}^{(0)}\right|^{2}}{2N_{0}^{2}}\frac{N_{0}^{2}k_{h}^{2}+f_{0}^{2}m^{2}}{N_{0}^{2}k_{h}^{2}}\;,$ (27)
$\hat{B}^{(0)}=\hat{\Theta}^{(0)}/\overline{\underline{\Theta}}$ is the leading-order buoyancy amplitude, and the horizontal wavenumber $k_{h}$ is defined as $k_{h}=\sqrt{k^{2}+l^{2}}$. The dispersion relation for internal gravity waves is given by
$\hat{\omega}^{2}=\frac{N_{0}^{2}k_{h}^{2}+f_{0}^{2}m^{2}}{\varepsilon^{4}k_{h}^{2}+m^{2}}\;.$ (28)
Therefore, we must express the leading-order tracer and wind amplitudes, $\hat{\Psi}^{(1)}$ and $\hat{\mathbf{V}}^{(0)}$, respectively, in terms of the buoyancy amplitude $\hat{B}^{(0)}$. The tracer amplitude is coupled to the wind amplitude via Equation˜18. The wind amplitude, in turn, is couple to the buoyancy amplitude through the polarization relations [4, 2]
$\hat{\mathbf{U}}^{(0)}$ $=\frac{i}{mN_{0}^{2}}\frac{\varepsilon^{4}\hat{\omega}^{2}-N_{0}^{2}}{\hat{\omega}^{2}-f_{0}^{2}}\left(\mathbf{k}_{h}\hat{\omega}-if_{0}\mathbf{e}_{z}\times\mathbf{k}_{h}\right)\hat{B}^{(0)}\;,$ (29)
$\hat{W}^{(0)}$ $=\frac{i\hat{\omega}}{N_{0}^{2}}\hat{B}^{(0)}\;,$ (30)
$\frac{c_{p}}{R}\underline{\overline{\Theta}}\hat{\Pi}^{(0)}$ $=\frac{i}{m}\frac{\varepsilon^{4}\hat{\omega}^{2}-N_{0}^{2}}{N_{0}^{2}}\hat{B}^{(0)}\;.$ (31)
@@
$M\mathbf{Z}^{(0)}=0\;,$ (32)
where the matrix $M$ is given by
$M=M(\mathbf{k},\hat{\omega})=\matrixquantity[-i\hat{\omega}&-f_{0}&0&0&ik\\ f_{0}&-i\hat{\omega}&0&0&il\\ 0&0&-i\hat{\omega}\varepsilon^{4}&-N_{0}&im\\ 0&0&N_{0}&-i\hat{\omega}&0\\ ik&il&im&0&0]\;,$ (33)
the components of the wave numbers are denoted by $\mathbf{k}=(k,l,m)^{\mathrm{T}}$, and the vector $\mathbf{Z}^{(0)}$ contains the wave amplitudes
$\mathbf{Z}^{(0)}=\left(\hat{U}^{(0)}\>,\>\hat{V}^{(0)}\>,\>\hat{W}^{(0)}\>,\>\hat{B}^{(0)}/N_{0}\>,\>\frac{c_{p}}{R}\overline{\underline{\Theta}}\hat{\Pi}^{(0)}\right)^{\mathrm{T}}\;.$ (34)
@@
By inserting Equations˜18, 29, 30, 28 and 27 into Equation˜25, we obtain an expression for $\mathcal{Q}^{(0)}$ in terms of the wave action density:
$\mathcal{Q}^{(0)}=-\frac{\varepsilon}{\overline{\rho}}\nabla_{\mathbf{X}}\cdot\left(\frac{f_{0}}{\hat{\omega}}\frac{m}{\varepsilon^{4}k_{h}^{2}+m^{2}}\mathcal{A}\mathbf{k}\times\nabla_{\mathbf{X}}\left\langle{\underline{\Psi}}\right\rangle\right)\;.$ (35)
@@ 2.3 Next-order gravity wave tracer flux convergence
As shown in the previous section, the leading-order tracer flux convergence $\mathcal{Q}^{(0)}$ arises from inertia-gravity waves and is the dominant gravity-wave forcing on tracers in rotating flows. Because the Coriolis frequency decreases toward the equator, this leading-order contribution weakens. In equatorial settings, such as those influenced by tropical convective gravity waves, the tracer forcing must come from higher-order terms. Although the global impact of the next-order convergence $\mathcal{Q}^{(1)}$ is likely small, it becomes relevant in idealized or weak-rotation regimes, including simulations without planetary rotation. For completeness and to enable accurate tests of parameterized waves under such conditions, we therefore derive a method to determine $\mathcal{Q}^{(1)}$ from the wave action density and wavenumbers.
The next-order flux convergence, $\mathcal{Q}^{(1)}$, is given in Equation˜26. Since $\mathcal{A}\propto\left|\hat{B}^{(0)}\right|^{2}$, both the leading-order and next-order tracer and wind amplitudes must be expressed in terms of the buoyancy amplitude. As described in the previous section, the leading-order amplitudes are coupled to $\hat{B}^{(0)}$ via Equations˜18, 29 and 30. The next-order tracer amplitudes can be obtained from the next-order wind amplitudes and the leading-order tracer amplitudes, as shown in Equation˜19. Therefore, the remaining task is to determine $\hat{\mathbf{V}}^{(1)}$ in terms of $\hat{B}^{(0)}$. Similar to the leading-order basic wave amplitudes, the next-order wind, buoyancy and Exner pressure amplitudes $\hat{\mathbf{V}}^{(1)}$, $\hat{B}^{(1)}=\hat{\Theta}^{(1)}/\overline{\underline{\Theta}}$, and $\hat{\Pi}^{(1)}$, respectively, satisfy a matrix equation
$M\mathbf{Z}^{(1)}=\mathbf{R}^{(1)}\;,$ (36)
where $M$ given in Equation˜33,
$\mathbf{Z}^{(1)}=\left(\hat{U}^{(1)}\>,\>\hat{V}^{(1)}\>,\>\hat{W}^{(1)}\>,\>\hat{B}^{(1)}/N_{0}\>,\>\frac{c_{p}}{R}\overline{\underline{\Theta}}\hat{\Pi}^{(1)}\right)^{\mathrm{T}}\;,$ (37)
@@
$\mathbf{R}^{(1)}=\left(R_{u}\>,\>R_{v}\>,\>R_{w}\>,\>R_{b}/N_{0}\>,\>R_{\pi}\right)^{\mathrm{T}}\;.$ (38)
The components of $\mathbf{R}^{(1)}$ are
$\begin{split}\mathbf{R}_{\mathbf{u}}^{(1)}=&-\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)\hat{\mathbf{U}}^{(0)}-\left(\hat{\mathbf{V}}^{(0)}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\mathbf{U}}}\right\rangle-\frac{c_{p}}{R}\left[\overline{\underline{\Theta}}\nabla_{\mathbf{X},h}\hat{\Pi}^{(0)}+\varepsilon\left(i\mathbf{k}_{h}\left\langle{\underline{\Theta}}\right\rangle\hat{\Pi}^{(0)}+\hat{\Theta}^{(0)}\nabla_{\mathbf{X},h}\left\langle{\underline{\Pi}}\right\rangle\right)\right]\;,\end{split}$ (39)
$\begin{split}R_{w}^{(1)}=&-\varepsilon^{4}\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)\hat{W}^{(0)}-\frac{c_{p}}{R}\left[\overline{\underline{\Theta}}\partial_{\mathrm{Z}}\hat{\Pi}^{(0)}+\varepsilon\left(im\left\langle{\underline{\Theta}}\right\rangle\hat{\Pi}^{(0)}+\hat{\Theta}^{(0)}\partial_{\mathrm{Z}}\left\langle{\underline{\Pi}}\right\rangle\right)\right]\;,\end{split}$ (40)
$\begin{split}R_{b}^{(1)}=&-\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)\hat{B}^{(0)}-\frac{\hat{\mathbf{V}}^{(0)}}{\overline{\underline{\Theta}}}\cdot\nabla_{\mathbf{X}}\left\langle{\underline{\Theta}}\right\rangle\;,\end{split}$ (41)
$\begin{split}R_{\pi}^{(1)}=&-\frac{1}{\overline{P}}\nabla_{\mathbf{X}}\cdot\left(\overline{P}\hat{\mathbf{V}}^{(0)}\right)\;,\end{split}$ (42)
with $\mathbf{R}_{\mathbf{u}}^{(1)}=\left(R_{u}^{(1)}\>,\>R_{v}^{(1)}\right)^{\mathrm{T}}$. These right-hand sides follow Achatz et al. [4]. However, in contrast to that work, we have included higher-order terms to facilitate a unified treatment of the two stratification cases (see Appendix˜A). Likewise, we retain the full expansion of the advecting large-scale winds. Using the polarization relations from Equations˜29, 30 and 31, all right-hand sides can be determined form the spatial and temporal dependence of the wavenumber and leading-order buoyancy amplitude $\hat{B}^{(0)}$. Once these terms are known, the next-order wave amplitudes $\mathbf{Z}^{(1)}$ can be obtained by solving Equation˜36.
@@
$\hat{B}^{(0)}=\left|\hat{B}^{(0)}\right|e^{i\varLambda}\;,$ (43)
where $\varLambda$ represents the large-scale wave phase. The absolute magnitude can be obtained from the wave energy $E_{w}$ via Equation˜27. While $E_{w}$ is directly available from the gravity wave model, the large-scale phase must be predicted from the prognostic equation
$\begin{split}\left(\partial_{\mathrm{T}}+\mathbf{c}_{g}\cdot\nabla_{\mathbf{X}}\right)\varLambda=&-\frac{f_{0}m}{2\left|\mathbf{k}\right|^{2}\hat{\omega}}\left(\mathbf{k}\times\nabla_{\mathbf{X}}\right)\left\langle{\underline{\mathbf{U}}}\right\rangle\\ &+\frac{|\mathbf{k}|^{2}}{2m\left(N_{0}^{2}-\varepsilon^{4}f_{0}^{2}\right)}\left[\frac{1}{\overline{\underline{\Theta}}}\left(\hat{\mathbf{c}}_{g}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Theta}}\right\rangle-\varepsilon N_{0}^{2}\frac{c_{p}}{R}\overline{\underline{\Theta}}\left(\hat{\mathbf{c}}_{g}\cdot\nabla_{\mathbf{X}}\right)\left\langle{\underline{\Pi}}\right\rangle\right]\\ &-\frac{f_{0}\hat{\omega}m^{2}}{k_{h}^{2}\left(N_{0}^{2}-\varepsilon^{4}f_{0}^{2}\right)}\left[\left(l\partial_{\mathrm{X}}-k\partial_{\mathrm{Y}}\right)\hat{\omega}+l\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)k-k\left(\partial_{\mathrm{T}}+\left\langle{\underline{\mathbf{U}}}\right\rangle\cdot\nabla_{\mathbf{X},h}\right)l\right]\end{split}$ (44)
where $\mathbf{c}_{g}=\nabla_{\mathbf{k}}\omega$ is the group velocity and $\hat{\mathbf{c}}_{g}=\nabla_{\mathbf{k}}\hat{\omega}$ is the intrinsic group velocity. We derive the prognostic equation by taking the imaginary part of the solvability condition, as outlined in Achatz et al. [3]:
$0=\mathbf{Z}^{(0)^{\dagger}}\mathbf{R}^{(1)}\;.$ (45)
In summary, $\hat{B}^{(0)}$ in Equation˜43 can be calculated using Equations˜27 and 44. The leading-order tracer and wind amplitudes can then be determined from Equation˜18 and Equations˜29 and 30, respectively. Next, by solving Equation˜36 for $\mathbf{Z}^{(1)}$ and using Equation˜19, the next-order tracer and wind amplitudes can be computed. Finally, the resulting $\hat{\Psi}^{(1)}$, can be substituted into Equation˜26 to obtain $\mathcal{Q}^{(1)}$.
@@ 3 Summary of equations in dimensional form
$\left[\left\langle{\underline{\mathbf{U}}}\right\rangle,\>\left\langle{\underline{W}}\right\rangle,\>\left\langle{\underline{\Psi}}\right\rangle,\>\left\langle{\underline{\Theta}}\right\rangle,\>\left\langle{\underline{\Pi}}\right\rangle\right]$ $\rightarrow\left[\frac{\left\langle{\mathbf{u}}\right\rangle}{\widetilde{U}},\>\frac{\left\langle{w}\right\rangle}{\varepsilon\widetilde{W}},\>\frac{\left\langle{\psi}\right\rangle}{\widetilde{\Psi}},\>\frac{\left\langle{\theta}\right\rangle}{\varepsilon^{2}T_{00}},\>\frac{\left\langle{\pi}\right\rangle}{\varepsilon^{2}}\right]\;.$ (52)
@@
$\hat{\omega}^{2}=\frac{N^{2}k_{h}^{2}+f^{2}m^{2}}{k_{h}^{2}+m^{2}}\;,$ (53)
@@
[…]
@@
$\mathbf{u}^{\prime}$ $=\frac{i}{mN^{2}}\frac{\hat{\omega}^{2}-N^{2}}{\hat{\omega}^{2}-f^{2}}\left(\mathbf{k}_{h}\hat{\omega}-if\mathbf{e}_{z}\times\mathbf{k}_{h}\right)b^{\prime}\;,$ (55)
$w^{\prime}$ $=\frac{i\hat{\omega}}{N^{2}}b^{\prime}\;,$ (56)
$c_{p}\overline{\theta}\pi^{\prime}$ $=\frac{i}{m}\frac{\hat{\omega}^{2}-N^{2}}{N^{2}}b^{\prime}\;.$ (57)
@@
$\psi^{\prime}$ $=-\frac{i}{\hat{\omega}}\mathbf{v}^{\prime}\cdot\nabla\left\langle{\psi}\right\rangle\;.$ (58)
@@
$\underbrace{\matrixquantity[-i\hat{\omega}&-f&0&0&ik\\ f&-i\hat{\omega}&0&0&il\\ 0&0&-i\hat{\omega}&-N&im\\ 0&0&N&-i\hat{\omega}&0\\ ik&il&im&0&0]}_{M}\underbrace{\matrixquantity[u^{\prime\prime}\\ v^{\prime\prime}\\ w^{\prime\prime}\\ b^{\prime\prime}/N\\ c_{p}\overline{\theta}\pi^{\prime\prime}]}_{\mathbf{Z}^{\prime\prime}}=\underbrace{\matrixquantity[R_{u}\\ R_{v}\\ R_{w}\\ R_{b}/N\\ R_{\pi}]}_{\mathbf{R}}\;,$ (59)
@@
$\left(\hat{\mathbf{U}}^{(1)}\>,\>\hat{W}^{(1)}\>,\>\hat{B}^{(1)}\>,\>\hat{\Pi}^{(1)}\>,\>\hat{\Psi}^{(2)}\right)\rightarrow\left(\frac{\mathbf{u}^{\prime\prime}}{\widetilde{U}}\>,\>\frac{w^{\prime\prime}}{\widetilde{W}}\>,\>\frac{\theta^{\prime\prime}}{\varepsilon^{2}\overline{\theta}}\>,\>\frac{\pi^{\prime\prime}}{\varepsilon^{3}}\>,\>\frac{\psi^{\prime\prime}}{\varepsilon\widetilde{\Psi}}\right)\frac{1}{\varepsilon}\;,$ (60)
and the components of $\mathbf{R}$ are
$\begin{split}\mathbf{R}_{\mathbf{u}}=&-\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)\mathbf{u}^{\prime}-\mathbf{v}^{\prime}\cdot\nabla\left\langle{\mathbf{u}}\right\rangle-c_{p}\left[\overline{\theta}\nabla_{h}\pi^{\prime}+i\mathbf{k}_{h}\left\langle{\theta}\right\rangle\pi^{\prime}+\theta^{\prime}\nabla_{h}\left\langle{\pi}\right\rangle\right]\;,\end{split}$ (61)
$\begin{split}R_{w}=&-\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)w^{\prime}-c_{p}\left[\overline{\theta}\partial_{z}\pi^{\prime}+im\left\langle{\theta}\right\rangle\pi^{\prime}+\theta^{\prime}\partial_{z}\left\langle{\pi}\right\rangle\right]\;,\end{split}$ (62)
$\begin{split}R_{b}=&-\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)b^{\prime}-\frac{\mathbf{v}^{\prime}}{\overline{\theta}}\cdot\nabla\left\langle{\theta}\right\rangle\;,\end{split}$ (63)
$\begin{split}R_{\pi}=&-\frac{1}{\overline{\rho}\overline{\theta}}\nabla\cdot\left(\overline{\rho}\overline{\theta}\mathbf{v}^{\prime}\right)\;.\end{split}$ (64)
@@
$i\hat{\omega}\psi^{\prime\prime}=-\left(\partial_{t}+\left\langle{\mathbf{u}}\right\rangle\cdot\nabla_{h}\right)\psi^{\prime}+\mathbf{v}^{\prime\prime}\cdot\nabla\left\langle{\psi}\right\rangle\;.$ (65)
Re-dimensionalized, the tracer flux convergence due to the Stokes drift $\mathcal{Q}^{(0)}$ is
$\mathcal{Q}^{(0)}=-\frac{1}{\overline{\rho}}\nabla\cdot\frac{f}{\hat{\omega}}\frac{m}{k_{h}^{2}+m^{2}}\mathcal{A}\mathbf{k}\times\nabla\left\langle{\psi}\right\rangle\;,$ (66)
representing the convergence of tracer fluxes due to the inertia-gravity wave Stokes drift from the gradient of the large-scale tracer. The next-order gravity wave-impact
$\begin{split}\mathcal{Q}^{(1)}=-\frac{1}{2\overline{\rho}\overline{\theta}}&\nabla\cdot\mathfrak{R}\left[\overline{\rho}\overline{\theta}\left(\mathbf{v}^{\prime}\psi^{\prime\prime*}+\mathbf{v}^{\prime\prime}\psi^{\prime*}\right)\right]\;,\end{split}$ (67)
arising from nonlinear effects associated with next-order wave fields $\mathbf{v}^{\prime\prime}$ and $\psi^{\prime\prime}$, generated by the first-order fields (Equations˜59 and 65). $\mathcal{Q}^{(1)}$ is determined through a sequential process, under the assumption that the wave energy $E_{w}$ is known: